f(x) is a 3rd Degree Polynomial Equation having Real coefficients, if it has 3, i and -i as Zeros (roots) Find a simple equation f(x) = 0 with these roots written in polynomial form (show work).
Answer by nerdybill(7384)

f(x) is a 3rd Degree Polynomial Equation having Real coefficients, if it has 3, i and -i as Zeros (roots) Find a simple equation f(x) = 0 with these roots written in polynomial form (show work).
.
(x-3)(x-i)(x+i) = 0
(x-3)(x^2+xi-xi-i) = 0
(x-3)(x^2-i) = 0
x^2(x-3)-i(x-3) = 0
x^3-3x^2-xi+3i = 0 (this is what they're looking for)